How do I calculate what volume of water should be added to 10.0 mL of 6.00 M HCL to produce a 0.500 M HCl solution?

1
Expert's answer
2021-09-07T02:09:05-0400

We have 10mL of 6.00 M HCL.


6.00M HCl means 6 moles of HCl are present in 1000 ml of its solution.


Thus , in 1ml of its solution 0.006 moles of HCl are present.


This implies , in 10ml 0.06 moles of HCl are present.


We want 0.5 M HCl which means we want 0.5 moles of it in 1000ml of its solution.


We have 0.06 moles of HCl .


Thus , volume of required to make it 0.5M is given by: moles of HCl / molarity of solution

"=\frac{0.06}{0.5} = 0.12 L"

Thus , volume of solution should be 0.12L = 120 ml.

Volume of Water to be added = total final volume - initial volume

= 120 - 10 = 110 ml